# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ VIII-B2 System Improvements
$\bullet$ Dynamic Resource Management Dynamic resource management optimizes performance and resource utilization in the fluctuating demands of edge-cloud environments through efficient, real-time resource allocation and reallocation.  Key aspects include algorithms that dynamically adjust resource allocation based on real-time workload characteristics and available resources [126], such as the energy-aware and trust-collaboration cross-domain resource allocation (ETCRA) algorithm, which minimizes system functions while adhering to latency and trust constraints [130].  Collaborative resource management coordinates resource allocation across multiple edge nodes and the cloud for enhanced system efficiency [15].  Additionally, integrating RL allows systems to learn optimal resource allocation strategies based on environmental feedback [21], while dynamic algorithms with initial allocation and real-time updating based on trust assessment improve performance [51]. Multi-timescale algorithms address the varying optimization periods of different sub-problems like service placement and task scheduling within edge-cloud cooperation [126].
@@
$\bullet$ Adaptive Optimization  Adaptive optimization dynamically adjusts system parameters and optimization strategies in response to real-time data and evolving conditions to maximize efficiency. For instance, the dynamic resource allocation mechanism in [21] ensures efficient scheduling of edge resources for global optimization.  Adaptive training with small batches, as proposed in Shoggoth [255], allows edge model adaptation under limited computing power.  Similarly, incremental learning approaches periodically retrain edge AI models using real-time data [267], optimizing prediction accuracy and reducing cloud dependency.  Dynamic model selection based on resources and application requirements [15] and decentralized adaptive scheduling using A3C learning and R2N2 [308] are other crucial aspects.  Additionally, the two-stage iterated greedy optimization (TIGO) algorithm addresses changing user requirements for dynamic microservice placement [120], while adaptive switching between fog and cloud instances based on response time enhances performance [309].
@@
$\bullet$ Fault Tolerance  Fault tolerance ensures reliable operation by implementing fault detection, isolation, and recovery strategies.  DeepFT [310], for example, uses a deep surrogate model for proactive fault prediction and optimized task scheduling.  Ensuring service resilience for critical nodes is addressed by confidence-aware resilience models like CAROL [311], which leverages generative neural networks for QoS prediction and system recovery.  Programmable asset orchestration at the edge contributes to operational resilience, with systems like RECS [312] dynamically managing services and resources during incidents.  Finally, distributed data auditing schemes, often incorporating techniques like FSS [188], ensure data integrity for reliable AI applications.
@@ IX Conclusion
In this survey, we comprehensively examine AI-driven edgecloud collaborative computing, focusing on distributed intelligence and model optimization. We first outline foundational architectures, then categorize key techniques into model optimization, resource management, and privacy-security enhancements. Additionally, we highlight trends in AI-driven resource allocation and model adaptation. Despite notable progress, challenges in system heterogeneity, security, and application diversity persist, urging further research into advanced AI techniques and emerging technologies like 6G, LLM integration, AI agents, embodied AI, edge quantum computing, etc.

### v5

@@ VIII-B2 System Improvements
$\bullet$ Dynamic Resource Management Dynamic resource management optimizes performance and resource utilization in the fluctuating demands of edge-cloud environments through efficient, real-time resource allocation and reallocation.  Key aspects include algorithms that dynamically adjust resource allocation based on real-time workload characteristics and available resources [fan2024collaborative], such as the energy-aware and trust-collaboration cross-domain resource allocation (ETCRA) algorithm, which minimizes system functions while adhering to latency and trust constraints [li2024energyaware].  Collaborative resource management coordinates resource allocation across multiple edge nodes and the cloud for enhanced system efficiency [tuli2023ai].  Additionally, integrating RL allows systems to learn optimal resource allocation strategies based on environmental feedback [xu2024fusion], while dynamic algorithms with initial allocation and real-time updating based on trust assessment improve performance [cao2020edge]. Multi-timescale algorithms address the varying optimization periods of different sub-problems like service placement and task scheduling within edge-cloud cooperation [fan2024collaborative].
@@
$\bullet$ Adaptive Optimization  Adaptive optimization dynamically adjusts system parameters and optimization strategies in response to real-time data and evolving conditions to maximize efficiency. For instance, the dynamic resource allocation mechanism in [xu2024fusion] ensures efficient scheduling of edge resources for global optimization.  Adaptive training with small batches, as proposed in Shoggoth [wang2023shoggoth], allows edge model adaptation under limited computing power.  Similarly, incremental learning approaches periodically retrain edge AI models using real-time data [joshi2024integration], optimizing prediction accuracy and reducing cloud dependency.  Dynamic model selection based on resources and application requirements [tuli2023ai] and decentralized adaptive scheduling using A3C learning and R2N2 [tuli2022dynamic] are other crucial aspects.  Additionally, the two-stage iterated greedy optimization (TIGO) algorithm addresses changing user requirements for dynamic microservice placement [he2024efficient], while adaptive switching between fog and cloud instances based on response time enhances performance [athanasopoulos2022ai].
@@
$\bullet$ Fault Tolerance  Fault tolerance ensures reliable operation by implementing fault detection, isolation, and recovery strategies.  DeepFT [tuli2023deepft], for example, uses a deep surrogate model for proactive fault prediction and optimized task scheduling.  Ensuring service resilience for critical nodes is addressed by confidence-aware resilience models like CAROL [tuli2022carol], which leverages generative neural networks for QoS prediction and system recovery.  Programmable asset orchestration at the edge contributes to operational resilience, with systems like RECS [moura2022resilience] dynamically managing services and resources during incidents.  Finally, distributed data auditing schemes, often incorporating techniques like FSS [li2024nextgeneration], ensure data integrity for reliable AI applications.
@@ IX Conclusion
In this survey, we presented a holistic analysis of AI-driven edge-cloud collaborative computing, emphasizing the foundational principle that “design-time” model optimization serves as a critical enabler for effective “run-time” resource management. Our work synthesizes key developments in model optimization, resource management, and security to construct a unified framework for the field, highlighting the synergistic relationship between creating efficient AI models and orchestrating their deployment across distributed infrastructure. By systematically connecting these domains, this review provides a comprehensive understanding of how to architect intelligent systems that are both powerful and resource-aware. By systematically structuring the co-design challenges and opportunities in model optimization and resource management, this survey extends beyond the scope of recent works and charts a clear path for future research in edge intelligence.
Despite substantial advancements, several open challenges persist that require focused research. Key issues include guaranteeing robust security and privacy in highly decentralized environments, achieving scalable and reliable performance with heterogeneous hardware, establishing standardized benchmarking protocols to ensure fair and reproducible evaluations, and promoting integrated hardware-software co-design for optimized performance. To address these gaps, future research should prioritize several high-impact directions, including: (1) Creating end-to-end confidential computing frameworks to protect data throughout the ECCC lifecycle; (2) Engineering unified orchestration platforms that streamline management and improve resource utilization across the computing continuum; (3) Establishing community-driven benchmarking suites to standardize evaluation and foster collaborative innovation; and (4) Investigating cross-layer optimizations that jointly consider AI models and emerging technologies like 6G, neuromorphic computing, and quantum edge computing to unlock new frontiers in distributed intelligence.